In a recent fundraiser, Daniela sold 5 less than twice as many coupon books as Andrew. Together, they sold a total of 28 coupon books. How many did Daniela sell?
Answer by CubeyThePenguin(3113)

Andrew = a
Daniela = 2a - 5

a + (2a - 5) = 28
3a - 5 = 28
3a = 33

a = 11

Daniela = 2a - 5 = 2(11) - 5 = 17